OpenStudy (anonymous):

The perimeter of a rectangle is 162 ft. The ratio of the length to the width is 5:4. Find the length of the rectangle. A.18 ft B.36 ft C.45 ft D.90 ft

OpenStudy (mathmale):

You could start by determining how the length relates to the width. If the ratio of the length to the width is 5:4, then the ratio length/width = l/w = 5/4. Solve for w in terms of l or for l in terms of w (doesn't matter which!). What next?

OpenStudy (anonymous):

So,you divide 5 and 4 to get 1.25.Since I did that what would I do next I honestly don't get this at all?...

OpenStudy (dumbcow):

 5x+5x+4x+4x = 162 18x = 162 x = 9 length = 5*9 = 45 flutter

OpenStudy (anonymous):

Oh I see that makes sense now Thank You!

OpenStudy (mathmale):

Please take a look at my original comment. The ratio of length (L) to width (W) is 5/4. Thus, L/W = 5/4. Cross multiplying, 4L = 5W. Should we choose to solve for L, L = 5W/4. Now substitute that back into the perimeter equation: 2W + 2L = 162 = 2W + 2(5W/4) Solve this for W. I get W = 36. Now calculate L: L = 5W/4 = ? Check your results by substituting L and W back into the original perimeter equation.

OpenStudy (mathmale):

dumbcow: you're no dumb cow. Your solution is quicker and more elegant than mine. Kudos.

OpenStudy (dumbcow):

@mathmale , i thank you but actually your solution by substitution is mathematically sound and just as quick if one is good at algebra. My solution is helpful for those who need a visual and may not be as adept at algebra